2 Ω 6A P2 10 Ω 5 Ω 4 Ω ΜΑ- 3Α
Answer #1

Givon circuit 22 un on V2 + + VI 101 vis 52 V2 & 40 v.13A Apply Nodal, al (vi) Node + V + VI-V2 + 6 = 0 V/5 2 V1 + V1 + 5V1 -Substituto oquation - 2v1 + 3 (12) = 36 -2 V1 = 36-36 Vi= 0 volts zole 52&or ac russ voltage is [bесото a short circuited) 22